find the moment of inertia of a uniform half-disc about an axis perpendicular to the plane and passing through its centre of mass. mass of this disc is M and radius is R.
 

Grade:11

1 Answers

Arun
6 years ago
mass of complete disc = M + M = 2M
the MOI of complete disc about the given axis is = 2MR2/2 = MR2
thus MOI of Half disc = MR2/2